# GUing: A Mobile GUI Search Engine using a Vision-Language Model

### v1

@@ 7.2. Vision Language Models for Mobile GUI
Vision-language models combines both the vision and language modalities, it has been trained on image-text pairs and can be applied to a range of tasks (Du et al., 2022; Zhang et al., 2023b).
Up until now, the main use case of VLMs in mobile GUI domain is the GUI captioning. GUI captioning refers to the task of generating a high-level summary of a screenshot to describe its contents and functionalities. The Screen2Words model (Wang et al., 2021), built on the foundations of the Transformer Encoder-Decoder architecture (Vaswani et al., 2017), uses multi-modal data, inclusive of the screenshot image, view hierarchy, and app description, to yield the screen caption. Contrarily, Spotlight, a vision-only approach employed for multiple tasks including widget captioning and screen captioning, possess architecture rooted in Vision Transformer (Dosovitskiy et al., 2021) and the T5 model (Raffel et al., 2020).
While some studies pertaining to GUI-related tasks may appear to employ VLMs, they do not in reality. For instance, XUI (Leiva et al., 2023) model does not address the language modality, it generates the screen caption with template-based natural language generation (NLG) engine. As presented on the previous section, GUIGLE (Bernal-Cardenas et al., 2019) and RaWi (Kolthoff et al., 2023) conducted text-based GUI retrieval by utilizing the textual content exhibited in the screenshots or the metadata, instead of employing the vision information for searching purposes. Such an approach may result in the omission of crucial vision information of the UI screen. To the best of our knowledge, no existing research employs VLMs for text-to-GUI retrieval.
@@ 7.3. Mobile GUI Datasets
7.3.1. Large Scale Dataset
The large-scale mobile GUI dataset comprises a vast collection of screenshots, thereby serving as a valuable resource for GUI retrieval.
@@ 7.3.1. Large Scale Dataset
Rico (Deka et al., 2017; Liu et al., 2018) is one of the largest GUI datasets in the literature. Deka et al. combined crowd sourcing and automation to mine design and interaction data from Android apps at runtime. It exposes visual, textual, structural, and interactive design properties of more than 66k unique UI screens from more than 9.3k Android apps. Several other dataset are built based on Rico (Leiva et al., 2020; Wang et al., 2021; Sunkara et al., 2022; Zaki and Abdallah, 2023). However, as the Rico dataset was created seven years ago, some of its UIs may be outdated for contemporary apps.
Chen et al. (Chen et al., 2020) developed an extensive dataset that comprises 54,987 screenshots derived from 7,748 Android apps, spanning 25 distinct app categories. These screenshots are obtained with automated GUI exploration (Chen et al., 2018).
Moran et al. (Moran et al., 2020) collected the ReDraw dataset in a fully automated manner by mining and executing the top-250 Android apps in each category on Google Play, excluding games. This yielded a total of 14,382 unique screenshots and 191,300 labeled GUI components after data cleaning.
WebUI (Wu et al., 2023b) is a large dataset of 400,000 rendered web pages associated with extracted metadata using automated web crawling. Although this dataset is not designed specifically for mobile apps, it can be useful for understanding mobile UI through transfer learning.
More and more automated GUI exploration methods are introduced to explore app UI automatically from iOS apps (Wu et al., 2023a) or Android apps (Chen et al., 2019a, 2023; Zhang et al., 2023a). Nonetheless, GUI datasets gathered through automated exploration may inadvertently omit important app features. Because the access to certain pages may necessitate app authorization or initial configurations, which are often absent in automatic exploration methods (Chen et al., 2023). Generally automated UI explorer are rather useful for UI testing than for getting design insights.
7.3.2. GUI Caption Dataset
A GUI caption dataset is essential for the training of vision language models (Radford et al., 2021; Singh et al., 2022; Li et al., 2022a, 2023).
Screen2Words dataset (Wang et al., 2021) is a subset of Rico, it contains 112,085 manually created descriptions across 22,417 unique screenshots. Moran et al. created Clarity dataset (Moran et al., 2022). They selected a subset of screenshots from the ReDraw dataset (Moran et al., 2020), and created a annotated dataset consisting of 45,998 descriptions for 10,204 screenshots from popular Android app. Instead of focusing on entire screens, Li et al. (Li et al., 2020) released a dataset containing 162,859 language phrases created by human workers for annotating 61,285 UI components across 21,750 unique UI screens.
8. Threats to Validity
This section discusses potential threats to the validity of our study.
Query Selection for Search Engine Evaluation. The selection of queries is crucial for evaluation, as the performance of a GUI search engine may vary depends on the query. For example, a GUI search engine that exhibits satisfactory results in the health and fitness domain may not necessarily produce the satisfactory outcomes in finance or education domains. During the evaluation process, it is imperative to consider the varying domains of applications. The study of Kolthoff et al. (Kolthoff et al., 2023) provided a gold standard of queries and corresponding GUIs. However, this dataset has already been exhausted for the training of RaWi, thereby rendering it unsuitable for testing within our study. Given the lack of prior queries, we instead sourced from articles published by companies that specialize in mobile app development services. The app features referenced in (Terekhov, 2023) span ten distinct domains, while (Hughart, 2023) details innovative features from even more domains. Based on this, we posit that these selected sets are representative enough for our evaluation.
@@ 8. Threats to Validity
Subjectivity of manual evaluation. As for every study that includes manual evaluation, the process of evaluation a search engine may introduce inherent bias. In an attempt to mitigate bias and subjectivity within the evaluation, we incorporated a number of strategic steps. The evaluation outcomes may be influenced by the evaluators’ preference towards specific search engines. To tackle this issue, we created an evaluation tool which mixed and shuffled the resulting screenshots originating from the three distinct search engines. As a result, the evaluator cannot discern the origin of a specific screenshot, thereby curbing potential bias among diverse search engines. Moreover, the evaluation process was carried out by four evaluators, each with a minimum of five years’ experience in software development. They conducted a careful evaluation based on a evaluation guide (e.g. stating what is or not a relevant GUI) and using a uniform evaluation tool. And each GUI was assessed by at least two evaluators to minimize bias attributed to individual preference.
@@ 9. Conclusion and Future Work
In this paper, we presented GUing, a GUI search engine, built upon UIClip, a novel vision-language foundation model for the mobile GUI domain. The UIClip model was trained using the GPSCap dataset, which is a comprehensive dataset comprised of 135k screenshot-caption pairs. In addition, we created the GPSRepo dataset to serve as the search source for GUing, encompassing a total of 303k screenshots. GPSCap and GPSRepo, were generated through a fully automatic pipeline, which enables to regularly update the datasets, e.g. with new screen designs. The pipeline enables the collection, classification, cropping, and caption recognition of app introduction images as found in Google Play as well as in other stores too. Our evaluation, including a benchmarking on various datasets and a manual assessment of search results, indicate that our approach outperforms the state-of-the-art methods for GUI retrieval. Moreover, the evaluation highlights the critical role that the GPSCap and GPSRepo datasets play in achieving the top performance of the vision-language models within GUI retrieval. There are several future directions to follow up on our work:
  • Increasing and diversifying the screenshot dataset. The performance of a vision-language model is profoundly correlated with the size of its training dataset. According to AppBrain (AppBrain, 2024b), Google Play includes approximately 2.36 million apps as of 2024. Our training dataset includes merely 117k apps. Mining additional app introduction images will thus likely improve the performance of GUing and relevance of the results. Moreover, other platforms like Apple iOS, Samsung Galaxy Watch, or Microsoft HoloLens have their stores with similar app introduction images. Learning and searching those screenshots can similarly inspire the design of corresponding apps.
  • Hybrid search methods. The inherent limitations of natural language can pose difficulties in describing the desired GUIs exhaustively through text. To enhance the retrieval process, it might be beneficial to support a more diverse range of search methods. For instance, sketch-to-GUI retrieval would enable users to find GUI designs (or related follw up screens) based on their sketched designs. Additionally, GUI-to-GUI retrieval would enable users to retrieve other GUIs that exhibit a resemblance to the original interfaces.
  • Evaluations with practitioners and integration into the app development workflows. Although our empirical evaluation show the efficacy of GUing, the reality of software development can be much more complex and nuanced. Thus, additional empirical studies with app developers, UI designers, and requirements engineers would likely lead to comprehensive feedback on real usage, scalability, further search parameters, as well as how GUing and UIClip can be ideally integrated into existing app development workflows and tools.

### v2

@@ 7.2. Vision Language Models for Mobile GUI
Vision-language models combines both the vision and language modalities. They are trained on image-text pairs and can be applied to a range of tasks (Du et al., 2022; Zhang et al., 2024). Until this work, the main use case of VLMs in the mobile GUI domain is the GUI captioning. This refers to the task of generating a high-level summary of a screenshot to describe its contents and functionalities. The Screen2Words model (Wang et al., 2021), built on the foundations of the Transformer Encoder-Decoder architecture (Vaswani et al., 2017), uses multi-modal data, inclusive of the screenshot image, view hierarchy, and app description, to yield the screen caption. Contrarily, Spotlight, a vision-only approach employed for multiple tasks including widget captioning and screen captioning, possess architecture rooted in Vision Transformer (Dosovitskiy et al., 2021) and the T5 model (Raffel et al., 2020).
While some studies pertaining to GUI-related tasks may appear to employ VLMs, they do not in reality. For instance, XUI (Leiva et al., 2023) model does not fully address the language modality. It generates the screen caption with template-based natural language generation (NLG) engine. As discussed above, GUIGLE (Bernal-Cardenas et al., 2019) and RaWi (Kolthoff et al., 2023) introduced text-based GUI retrieval approaches by using the textual content exhibited in the screenshots or the metadata, instead of employing the vision information for the search. This may result in the omission of crucial vision information of the UI screen. To the best of our knowledge, no existing research employs VLMs for text-to-GUI retrieval.
Recently, two related VLMs, UIClip (Wu et al., 2024) and Ferret-UI (You et al., 2024), have been presented in parallel to our work. UIClip, like our model GUIClip, is a fine-tuned version the CLIP model. However, their purposes differ significantly. While GUIClip is designed for GUI retrieval (as app vendors describe what is implemented on the corresponding screen), UIClip is trained to assess the design quality and visual relevance of a UI. Ferret-UI, on the other hand, is a multimodal foundation model that goes beyond VLMs by supporting more input formats, such as points, boxes, and scribbles on screenshots. This model is capable of performing various tasks (e.g., widget classification, icon recognition, OCR) with flexible input formats, as well as grounding tasks (e.g., finding widgets, icons, or text, and listing widgets on a screen).
@@ 7.3. Mobile GUI Datasets
Large-scale mobile GUI datasets comprise a vast collection of screenshots and serve as a valuable resource for GUI retrieval. Rico (Deka et al., 2017; Liu et al., 2018) is one of the largest datasets in the literature. Deka et al. (Deka et al., 2017) combined crowdsourcing and automation to mine design and interaction data from Android apps at runtime. Their dataset exposes visual, textual, structural, and interactive design properties of more than 66k unique UI screens appx. 9.3k Android apps. Several other datasets (Leiva et al., 2020; Sunkara et al., 2022; Zaki and Abdallah, 2023) are built on Rico, including the Screen2Words dataset (Wang et al., 2021) used in this paper. As Rico was created seven years ago, some of its UIs might be outdated for contemporary apps as our evaluation suggests.
Chen et al. (Chen et al., 2020) developed an extensive dataset that comprises almost 55k screenshots from 7,748 Android apps. The screenshots were obtained through automated GUI exploration (Chen et al., 2018). Moran et al. (Moran et al., 2020) collected the ReDraw dataset in a fully automated manner by mining and executing the top-250 Android apps in each category on Google Play, excluding games. This yielded a total of 14,382 unique screenshots and 191k labelled GUI components. Clarity (Moran et al., 2022) is an annotated subset of ReDraw consisting of 45,998 descriptions for 10,204 screenshots of popular apps. The descriptions were created by crowd workers using Amazon Mechanical Turk. Instead of focusing on entire screens, Li et al. (Li et al., 2020) released a dataset containing appx. 162,859 phrases created by human workers for annotating 61k UI components across 21,750 unique screens.
Generally, more and more crowdsourced or automated GUI exploration methods are used to (semi)automatically explore UIs of iOS apps (Wu et al., 2023) or Android apps (Chen et al., 2019a, 2023; Zhang et al., 2023). While certainly efficient and scalable, automated UI exploration rather targets UI testing than GUI design inspiration. GUI datasets gathered through crowdsourced or automated exploration may omit important app features sine the access to certain UIs may necessitate app authorisation or initial configurations (Chen et al., 2023). In addition, these datasets are not curated by actual app designers and vendors. Recent studies showed that app vendors and app users (or the crowd) often use different vocabularies to describe app features (Haering et al., 2021a, b). The dataset introduced in our work ScreenRepo and SCapRepo aims at closing this gap. While we also used automated techniques for the data collection and cleaning, the screen-caption pairs in SCapRep are curated by various app vendors, describing what a screen is supposed to do and what problem for the user does it solve.
8. Discussion
We discuss the implications of our work for research and practice together with potential threats to validity.
8.1. Implications for Software Practitioners
Our work has several potential implications for software practitioners, particularly for app designers and requirements engineers. Our search engine primarily enables app designers and developers to get inspirations on how their screens can look like, for instance in an early project phase, when only a list of app features is available. Particularly in small development teams, which often lack resources and expertise for GUI design and usability engineering (Bornoe and Stage, 2013), this can provide a substantial support. Despite the general trend with Generative AI for more code generation and automation, we think that it remains crucial in professional app development to have the designer (or developer) in the loop (Wei et al., 2024). GUing is such a tool that uses AI to speed-up design tasks while ensuring the crucial role of human creativity.
GUing can also serve as a rapid prototyping tool. It can, e.g., be used during requirements interviews and in workshops, to retrieve screenshots based on identified requirements, speeding up their refinement and validation. This by itself, increases stakeholder engagement and helps mitigate development risks. Additionally, retrieved screenshots serve as a visual aid for illustrating and documenting the requirements and reducing potential ambiguity (Montgomery et al., 2022).
Given the significant impact of GUI on user experience and retention (Chen et al., 2021), and as GUI designs and user preferences evolve rapidly, it is essential for app designers to continuously observe and learn from trends and best practices. GUing leverages app introduction images curated by actual domain and app vendors. Our pipeline for image processing and model training allows for a seamless addition of new images to the GUing repository, ensuring a high quality and up-to-dateness. This helps designers learn the latest UI design trends. When augmented with download, rating, and feedback trends, retrieved screens can help not only get inspiration but also estimate user trends (Martens and Maalej, 2019b).
8.2. Implication for Software Researchers
There has been extensive research on app store mining over the last decade, with the majority of studies focusing on app reviews and descriptions (Al-Subaihin et al., 2021; Martin et al., 2017; Maalej et al., 2024). Less work have delved into the analysis of other developer information shared in the stores, such as app introduction images (Chen et al., 2019b; Feng et al., 2022). Our work highlights the potential of this information, particularly as it is curated, usually exhibits a high quality (as created by professionals), and captures various modalities (image, text, videos, etc.). We think that researchers should focus more on investigating this information and combining it with crowd-data for building multi-modal foundation models for software engineering.
From the app introduction images, we developed and shared the SCapRepo dataset and subsequently the vision-language model GUIClip. The evaluation results confirms the power for these models to learn and combine modalities, not only for screenshot retrieval but also at least for GUI classification and sketch-to-GUI retrieval. We think that GUIClip and similar foundation models tuned on software design data have the potential to be adapted for other GUI-related tasks, including GUI captioning and GUI generation. As a foundational model, CLIP model itself is the basis to numerous models (Zhang et al., 2024), such as the image captioning model ClipCap (Mokady et al., 2021) and the image generation model Stable Diffusion (Rombach et al., 2022). By replacing the CLIP model with tuned, domain specific models such as GUIClip, it is possible to enhance these derivative models’ capabilities in GUI captioning and GUI generation. Researchers should investigate this in details, not only with regard to model performance (or accuracy) but also with regard to quality of output, risk of prediction faults, as well as design and engineering workflows to be adapted (Wei et al., 2024).
8.3. Threats to Validity
This section discusses potential threats to the validity of our work.
8.3.1. Internal and Construct Validly
@@
As for every study that includes manual evaluation, the process of evaluating the usefulness of the search results may have included inherent biases. Particularly, the evaluation outcomes may have been influenced by the evaluators’ preferences towards specific search engines. In an attempt to mitigate this potential observer bias, we incorporated a number of steps. We created an evaluation tool which mixes and shuffles the resulting screenshots originating from the three distinct search engines. As a result, the evaluator cannot discern the origin of a specific screenshot, thereby curbing potential bias among diverse search engines. Moreover, the evaluation process was carried out by four evaluators, each with a minimum of five years’ experience in software development. They conducted a careful evaluation based on a evaluation guide (e.g. stating what is a relevant GUI and what is not) and using a uniform evaluation tool. Finally, each query was independently assessed by at least two evaluators to minimise potential mistakes.
8.3.2. External Validity
The selection of queries is crucial for the evaluation, since the search performance may vary depending on the query. For example, a GUI search engine that exhibits satisfactory results for the health and fitness domain may not necessarily produce the same outcome for the finance or education domains. It is thus imperative to consider varying domains and apps in the evaluation. The study of Kolthoff et al. (Kolthoff et al., 2023) provides a gold standard of queries and corresponding GUIs. However, this dataset has already been exhausted for the training of RaWi, which makes it unsuitable for our comparative study. Given the lack of alternative query datasets, we searched for articles published by companies specialised in mobile app development. This decision could potentially introduce a selection bias. The app features extracted from the articles span ten distinct domains (Terekhov, 2023), plus innovative features from multiple additional domains (Hughart, 2023). Therefore, we think that these sets are representative enough for the purpose of our evaluation. Nevertheless, reproducing our study with additional queries, scenarios, and evaluators would certainly further strengthen the generalisablity of its results.
@@ 9. Conclusion and Future Work
In this paper, we presented GUing, a GUI search engine, built upon GUIClip, a novel vision-language foundation model for the domain of mobile apps. The GUIClip model was trained using SCapRepo, a comprehensive dataset comprised of 135k screenshot-caption pairs created by actual app experts. In addition, we presented ScreenRepo, a dataset that serves as a large repository for GUing. SCapRepo and ScreenRepo were created through an automated pipeline, enabling an easy update to the datasets, e.g. when new screen designs emerge.
Our evaluation, including a benchmarking on various datasets and a manual assessment of search results, indicates that GUing outperforms state-of-the-art approaches for GUI retrieval. Moreover, the evaluation highlights the critical role that the SCapRepo and ScreenRepo datasets play in achieving the top performance of our vision-language models: not only for GUI retrieval but also for GUI classification and sketch-to-GUI retrieval in the mobile app domain. From here, there are several follow-up future directions:
  • Increasing and diversifying the screenshot dataset. The performance of a vision-language model is correlated with the size of its training data. According to Statista (Statista, 2024), Google Play includes approximately 2.43 million apps as of 2023. Our training dataset includes merely 117k apps. Mining additional app introduction images will thus likely improve the performance of GUing and relevance of the results. Moreover, other platforms like Apple iOS, Samsung Galaxy Watch, or Microsoft HoloLens also have stores with similar app introduction images. Learning and searching those screenshots can improve the engine and inspire the design of corresponding apps.
  • Hybrid search methods. The inherent limitations of natural language can pose difficulties in precisely describing the desired GUIs through text. To enhance the retrieval process, it might be beneficial to support a more diverse range of search methods and data. For instance, sketch-to-GUI retrieval would enable users to find GUI designs (or related follow up screens). Additionally, GUI-to-GUI retrieval would enable users to retrieve other GUIs that exhibit a resemblance (or other dependencies) to the original interfaces.
  • Evaluations with practitioners and integration into the development workflows. Although our empirical evaluation show the efficacy of GUing, the reality of software development can be more complex and nuanced. Thus, additional empirical studies with app developers, UI designers, and requirements engineers would likely lead to comprehensive feedback on real usage, scalability, further search parameters, as well as how GUing and GUIClip can be ideally integrated into existing app development workflows and tools (Maalej, 2009).